OpenStudy (anonymous):

A Fair eight-sided die with faces numbered 1 through 8 is rolled four times. What is the probability that exactly one 3 is rolled, and that 3 is rolled on the 3rd roll? Express your answer as a common fraction

OpenStudy (anonymous):

exactly one 3 is \[4\times (\frac{1}{8})(\frac{7}{8})^3\]

OpenStudy (anonymous):

exactly one three on the third roll is \[(\frac{1}{8})\times (\frac{7}{8})^3\]

OpenStudy (anonymous):

is that (78)3\[\left(\begin{matrix}7 \\ 8\end{matrix}\right)\] to the power of the 3? or is it to the power of 3 on both